OpenStudy (anonymous):

Help Please, I am trying to help my mom with some homework and I want to make sure the problems I am not sure about is right, thank you.

OpenStudy (anonymous):

\[\frac{ 7x+35 }{ x ^{2}+5x}\]

OpenStudy (anonymous):

\[\frac{ x+5 }{ x ^{2}-4x-45 }\]

OpenStudy (whpalmer4):

are you asked to simplify these expressions?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank yall so much.

OpenStudy (whpalmer4):

What do you have for answers? We'll check them for you.

OpenStudy (tkhunny):

FACTOR, FACTOR, FACTOR!!! \(x^{2} + 5x = x(x+5)\) \(7x+35 = 7(x+5)\) \(x^{2} - 4x - 45 = (What)(What)\)

OpenStudy (anonymous):

whpakmer4 I have for the 1st on is 7/x and for the 2nd I had 1/3x-9

OpenStudy (whpalmer4):

First one is correct.

OpenStudy (whpalmer4):

second one is not.

OpenStudy (anonymous):

okay um...

OpenStudy (whpalmer4):

show me your factoring of the denominator of the second problem

OpenStudy (anonymous):

I a sorry brb

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

\[\frac{ x+5 }{ x ^{2} -4x-45}\] \[\frac{5}{x-4x-45}\]

OpenStudy (anonymous):

then I simplified 5/-45

OpenStudy (tkhunny):

No. Absolutely not. Factor the denominator. Look at those "What's" I gave you up above.

OpenStudy (anonymous):

-9 and 5

OpenStudy (whpalmer4):

not -9 and 5, \((x-9)\) and \((x+5)\)

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

so you've got \[\frac{x+5}{(x+5)(x-9)}\]

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, can you simplify that any more?